OpenStudy (anonymous):

5x^2-5x-30=0 (Split the middle term)

OpenStudy (akashdeepdeb):

In these kinds of questions \[Ax^2 + Bx + C = 0\] You have to find 2 numbers that would MULTIPLY to give AC And that will add or subtract to give B In this case \[5x^2-5x-30=0\] 5x can be split as 15x - 10x Because 15 * 10 give us AC = 150 And 15-10 give us 5 that is B Getting me?

OpenStudy (anonymous):

Yeah.Thank you :)

OpenStudy (akashdeepdeb):

:)